QUESTION 1 Change in Consumption Change in Income Change in Saving Assumed Increase in Investment Second Round All Other Rounds Totals $20 $4.00 $12.80 $51.20 $20.00 Refer to the given table, which illustrates the multiplier process. The marginal propensity to consume is O 0.75 0.8 0.5 0.9
Answer #1

Answer is option B) MPC = .8

as ∆Y = ∆ C + ∆ S

& Multiplier is , m = 1/ MPS

MPS = ∆S/∆Y

= 4/20 = .2

Thus as MPS + MPC = 1

So MPC = 1-MPS = .8